### v1

@@ 5.1 Data Requirements and Computational Intensity
Upon investigating the deployment of ViTs within the realms of DA and DG methods, our research has identified a range of challenges that merit closer examination. ViTs have emerged as a pivotal enhancement in computer vision capabilities, promising significant advances across various applications. However, their adoption in real-world scenarios is fraught by considerable challenges, primarily due to their extensive data requirements. The necessity for large-scale datasets, such as JFT-300M, for effective training, underscores this challenge, as [16] have pointed out. This substantial reliance on voluminous, pre-trained models, emphasized by [17], necessitates access to high-quality data, a critical aspect particularly when training ViTs from scratch on more constrained datasets, a scenario highlighted by [29]. This dependency presents a significant barrier, wherein the effectiveness of ViTs, upon adaptation through transfer learning and fine-tuning, is closely linked to the quality of pre-existing object detectors. This necessitates a meticulous application of these methodologies, as evidenced by the works of [21] and [157]. In addition, a big challenge is when the models need to learn from a very small number of examples, a method known as few-shot learning. Situations that require high safety standards or have very limited data, show how hard it can be to adjust ViTs under different conditions. Recently, new methods like source-free domain adaptation and test time adaptation have shown promise in making models more generalized. Even though researchers made progress in overcoming biases from the source data, they still have a lot to learn about how to manage uncertainty when making these adjustments. This area presents significant opportunities for further research. Finally, the inherent self-attention mechanisms of ViTs, especially in models with a vast number of trainable parameters, introduce a layer of complexity that demands substantial computational resources. This further complicates their deployment in practical scenarios, as discussed by [61].
@@ 5.4 Uncertainty-Aware Vision Transformers
In our analysis of ViTs and their proficiency in navigating distribution shifts, we’ve highlighted their potential to enhance model generalization through various techniques. A particularly promising, yet under explored approach is integrating uncertainty quantification methods with ViTs. This integration enables models to provide predictions along with confidence levels, making decision-making more transparent. The presence of uncertainties, amplified by distribution shifts, is not merely an additional challenge but a crucial aspect of real-world environments’ unpredictability. Employing uncertainty-aware ViTs to detect and improve model generalizability presents a significant research opportunity. Future studies should delve into how uncertainties influence the adaptation and generalization capabilities of ViTs, emphasizing the integration of uncertainty quantification methods. Such investigative efforts are crucial for gaining a thorough understanding of how modern vision networks can exploit uncertainties to enhance the field of domain adaptation and generalization.
This review stands as the first to comprehensively gather most of the recent work on utilizing ViTs to address distribution shifts within the framework of DA and DG approaches. The increasing number of publications on this topic showcases a growing interest, underscoring the rapid evolution of the field and the keenness among researchers to explore further the capabilities of ViTs. As researchers, we see a promising future in using ViTs to address distribution shifts. This survey aims to offer a perspective on the evolution of ViTs, guiding future research toward the creation of more robust and versatile deep learning models.

### v2

@@ 5.1 Data Requirements and Computational Intensity
Upon investigating the deployment of ViTs within the realms of DA and DG methods, our research has identified a range of challenges that merit closer examination. ViTs have emerged as a pivotal enhancement in computer vision capabilities, promising significant advances across various applications. However, their adoption in real-world scenarios is fraught by considerable challenges, primarily due to their extensive data requirements. The necessity for large-scale datasets, such as JFT-300M, for effective training, underscores this challenge, as [17] have pointed out. This substantial reliance on voluminous, pre-trained models, emphasized by [18], necessitates access to high-quality data, a critical aspect particularly when training ViTs from scratch on more constrained datasets, a scenario highlighted by [29], and recently discussed in [164]. This dependency presents a significant barrier, wherein the effectiveness of ViTs, upon adaptation through transfer learning and fine-tuning, is closely linked to the quality of pre-existing object detectors. This necessitates a meticulous application of these methodologies, as evidenced by the works of [22], [165] and scalable approaches like [164]. In addition, a big challenge is when the models need to learn from a very small number of examples [166], a method known as few-shot learning. Situations that require high safety standards or have very limited data, show how hard it can be to adjust ViTs under different conditions. Recently, new methods like source-free domain adaptation and test time adaptation have shown promise in making models more generalized. Even though researchers made progress in overcoming biases from the source data [166], they still have a lot to learn about how to manage uncertainty when making these adjustments. This area presents significant opportunities for further research. Finally, the inherent self-attention mechanisms of ViTs, especially in models with a vast number of trainable parameters, introduce a layer of complexity that demands substantial computational resources. This further complicates their deployment in practical scenarios, as discussed by [63].
@@ 5.4 Uncertainty-Aware Vision Transformers
In our analysis of ViTs and their proficiency in navigating distribution shifts, we’ve highlighted their potential to enhance model generalization through various techniques. A particularly promising, yet under explored approach is integrating uncertainty quantification methods with ViTs [170]. This integration enables models to provide predictions along with confidence levels, making decision-making more transparent. The presence of uncertainties, amplified by distribution shifts, is not merely an additional challenge but a crucial aspect of real-world environments’ unpredictability. Employing uncertainty-aware ViTs to detect and improve model generalizability presents a significant research opportunity. Future studies should delve into how uncertainties influence the adaptation and generalization capabilities of ViTs, emphasizing the integration of uncertainty quantification methods. Such investigative efforts are crucial for gaining a thorough understanding of how modern vision networks can exploit uncertainties to enhance the field of domain adaptation and generalization.
This review is the first to comprehensively gather recent work on using ViTs to address distribution shifts in DA and DG approaches. The growing number of publications highlights increasing interest and rapid evolution in the field. We see a promising future for ViTs in addressing distribution shifts and aim to guide future research toward creating more robust and versatile deep learning models.
Data availability The data that support the findings of this study are publicly available and will be provided upon request.